### v3

# DEEP CLUSTERING USING ADVERSARIAL NET BASED CLUSTERING LOSS

Abstract
Deep clustering is a recent deep learning technique which combines deep learning with traditional unsupervised clustering. At the heart of deep clustering is a loss function which penalizes samples for being an outlier from their ground truth cluster centers in the latent space. The probabilistic variant of deep clustering reformulates the loss using $KL$ divergence. Often, the main constraint of deep clustering is the necessity of a closed form loss function to make backpropagation tractable. Inspired by deep clustering and adversarial net, we reformulate deep clustering as an adversarial net over traditional closed form $KL$ divergence. Training deep clustering becomes a task of minimizing the encoder and maximizing the discriminator. At optimality, this method theoretically approaches the $JS$ divergence between the distribution assumption of the encoder and the discriminator. We demonstrated the performance of our proposed method on several well cited datasets such as MNIST, REUTERS and CIFAR10, achieving on-par or better performance with some of the state-of-the-art deep clustering methods.
1 Introduction
Deep neural network such as the autoencoder is applied to many signal processing domains such as speech and audio processing [1, 2], image clustering [3, 4] and medical data processing [5, 6]. While the latent space of autoencoder is well suited for dimension reduction through reconstruction loss [7], the latter is not optimized for clustering/ classification since class labels cannot be used in the reconstruction loss [8]. A recent autoencoder technique known as the Euclidean distance based clustering (ABC) [7], utilize class label information in neural network by introducing a loss function known as the deep clustering loss. The goal is to minimizes the Euclidean distance in the latent space between samples and partitioning learnt by clustering algorithms e.g. Kmeans or Gaussian mixture model (GMM). In other words, the latent space of an autoencoder learnt using reconstruction loss will look different when we further perform K-means on the latent space. A recent probabilistic approach to ABC uses KL divergence (KLD) [4, 9] and assumes that both the variational autoencoder (VAE) [10] latent space and clustering approach are Gaussian distributed. We can compare the similarities and differences between deep clustering and VAE to better understand the former. The neural network of VAE and deep clustering are optimized by backpropagating samples in the latent space to update the encoder weights. Also, both deep clustering and VAE share a similar goal of modeling each input image as a sample draw of the Gaussian distribution in the latent space. However, their Gaussian distributions are different. VAE enforce samples in the latent space to be closely distributed by a Gaussian distribution with continuous mean $\mu$ and variance $\sigma$ i.e. $z\sim\mu+\sigma\cdot\mathcal{N}(0,1)$. Whereas the latent space of deep clustering approaches a Gaussian mixture model (GMM) and each Gaussian has a set of fixed parameters i.e. $z\mathcal{\sim N}(\eta_{k},\tau_{k})$, $\forall k\in K$. The goal of VAE is to enforce all class samples in the latent space to be distributed by continuous mean and variance. The goal of deep clustering is to enforce class samples in the latent space to be distributed by a discrete set of mean and variance i.e. $N(\eta^{*},\tau^{*})$. Thus, only deep clustering is associated with clustering, making it more suitable for unsupervised classification. Deep clustering is not without problem. Mainly, when backpropagating from a neural network, we require a closed form loss function for tractability [11]. Most deep learning methods including VAE restrict to the assumption of a KLD between two Gaussians to obtain a closed form loss function for backpropagating. When generalizing to $f$-divergence between two Gaussians, closed form solution does not exist. A well known workaround to this problem is to approximate the JS divergence (JSD) between two Gaussians using adversarial net. In fact, adversarial net is quite versatile as shown in Table 1. For different applications, the discriminator can be seen as performing a specific task different from the discriminators in other adversarial net. We distinguish the proposed work from other “adversarial net based X” in Table I where X can be any discriminator task. In GAN, X refers to image generation. Other instances of “adversarial net based X” includes AAE, DASC and etc. In our approach, “Deep clustering using adversarial net based clustering loss” (DCAN), X refers to clustering. More specifically, in DCAN the JSD between latent space and clustering is approximated by adversarial net. More details are discussed in the next section.
Adversarial net based X Discriminator
GAN [11] $\left\{\begin{array}[]{c}if\;x\sim p(data),\;T=1\\ else,\;T=0\end{array}\right.$ $x$ is from the dataset?
DASC [12] $\left\{\begin{array}[]{c}if\;z\sim p(\varsigma^{*}),\;T=1\\ else,\;T=0\end{array}\right.$ $z$ is from an assigned subspace?
AAE [13] $\left\{\begin{array}[]{c}if\;z\sim\mu+\sigma\cdot\mathcal{N}(0,1),\;T=1\\ else,\;T=0\end{array}\right.$ $z$ is from a Gaussian prior?
DCAN (proposed) $\left\{\begin{array}[]{c}if\;z\sim p(z\mid\theta^{*}),\;T=1\\ else,\;T=0\end{array}\right.$ $z$ is from an assigned cluster?
Table 1: Different strategies of using adversarial net based X
1.1 Adversarial net based X
The optimization of adversarial net [11] in general centers around minimizing the encoder/generator and maximizing the discriminator. However, as mentioned “adversarial net based X” are versatile and we highlight their similarities and differences as follows: In GAN, the generator is responsible for generating an image when given a sample from the latent space. The discriminator of GAN checks whether the generated image belongs to the distribution of the original dataset. Similarly, in DCAN we minimize the encoder and maximize the discriminator but for different purpose. In DCAN, the encoder receives an image and outputs a sample in the latent space. GAN trains the generator to be sensitive to the variance of the image class i.e. small displacement of the sample in the latent space can result in huge image variance. DCAN trains the encoder to be insensitive to the class variance i.e. different images from the same class populate tightly close to the cluster center in the latent space. Both DCAN and AAE [13] utilize the adversarial net in totally different ways despite having similar architecture (encoder acting as generator). The goal of AAE is identical to VAE i.e. the AAE encoder models a sample in the latent space as a sum and product of $\mathcal{N}(0,1)$. On the other hand, the goal of DCAN is not identical to VAE. DCAN models a sample in the latent space, as the sample draw from a cluster (e.g. Kmeans or Gaussian mixture model), specifically the sample draw of a $1-of-K$ cluster. The goal of DASC [12] is subspace clustering, and it is different from deep clustering. Like principal component analysis, DASC perform unsupervised learning by decomposing the dataset into different individual eigenvectors, each eigenvector as orthogonal to each other as possible. Samples from each classes should ideally reside on their respective $K^{th}$ subspaces, just as each sample should ideally fall within their respective $K^{th}$ clusters in deep clustering. In DASC, the latent space of DASC is trained to behave like subspace clustering. Thus the adversarial net of all four methods, GAN (models the dataset), AAE (models VAE loss), DASC (models subspace clustering) and DCAN (models deep clustering) are optimized in totally different ways as summarized in Table I.
2 Proposed method
Figure 1: Proposed deep clustering loss using adversarial net. It uses the discriminator and encoder architecture inspired by AAE.
One motivation for using a JSD approach is that it does not suffer from asymmetry unlike KLD [14, 11]. However, there is no closed form solution available for the JSD between two data distributions. GAN [11] overcome the need for a closed form solution by employing an adversarial training procedure. GAN approximates JSD or $D_{JS}\left(p_{data}\parallel p_{g}\right)=E_{x\sim p_{data}}\left[\log D_{G}(x)\right]+E_{x\sim p_{g}}\left[\log\left(1-D_{G}\left\{x\right\}\right)\right]$ at optimality, where $p_{data}$ is the real data distribution and $p_{g}$ is the generated data distribution. In the JSD between two Gaussian distributions for deep clustering, we seek $D_{JS}\left(\>p(z\mid\theta^{*})\parallel q(z\mid x)\>\right)=E_{z\sim p(z\mid\theta^{*})}\left[\ln D(z)\right]+E_{z\sim q(z\mid x)}\left[\ln\left(1-D\left\{z\right\}\right)\right]$, whereby $z\sim p(z\mid\theta^{*})$ refers to the sample distributed from clustering and $z\sim q(z\mid x)$ refers to the sample distributed from the encoder. Unlike the discriminator of GAN in the original space, the discriminator of DCAN works in the latent space. We can visualize $D_{JS}\left(\>p(z\mid\theta^{*})\parallel q(z\mid x)\>\right)$ using Fig. 1. The discriminator tries to distinguish samples from clustering, $z\sim p(z\mid\theta^{*})$ versus samples generated from the encoder, $z\sim q(z\mid x)$.
2.1 Training DCAN
We train DCAN by feeding it with $1...N$ positive and negative samples. We refer to each $n^{th}$ sample of $x$ as $x^{\left(n\right)}$ with dimension $I$. In the forward pass in Fig 1, samples enter the discriminator through the encoder and clustering. Samples generated from clustering1 are positive i.e. $T=1$ and samples generated from encoder are negative i.e. $T=0$. We define the neural network output and target label as $y=\left\{0,1\right\}$ and $T=\left\{0,1\right\}$ respectively. Both $y$ and $T$ have the same dimension $d$. We update the discriminator weight using DCAN loss in eqn (1). Since the samples in the latent space displaces each time the encoder weight changes, clustering parameters $\theta_{k}=\{\eta_{k},\tau_{k},\pi^{(n)}\forall k\}$ has to be re-computed each time the discriminator weight is updated. The encoder $q(z\mid x)$ maps $x$ to $z$ in the latent space. However, we desire to minimize between $q(z\mid x)$ and $p(z\mid\theta^{*}).$ Meaning that we desire to produce a sample $z$ from $q(z\mid x)$ that is as close as possible to the sampling from $p(z\mid\theta^{*})$. Similarly, DCAN loss in eqn (1) is used to update the encoder weight.
$\begin{array}[]{c}L_{DCAN}=\left\{\frac{1}{N}\sum_{n=1}^{N}\;\underset{x^{\left(n\right)}\sim p_{data}}{E}\left[\ln\left(1-D\left(G(x^{\left(n\right)})\right)\right)\right]\right\}_{T=0}\\ +\left\{\frac{1}{N}\sum_{n=1}^{N}\;\underset{z^{\left(n\right)}\sim p(z|\theta^{*})}{E}\left[\ln D\left(z^{\left(n\right)}\right)\right]\right\}_{T=1}\end{array}$ (1)
3 Experiments
We selected some benchmarked datasets used by deep clustering for our experiments in Tables 2. There are several factors affecting deep clustering. i) The clustering algorithm used in the latent space e.g. Kmeans or GMM. ii) The number of hidden layers and nodes of the neural network. iii) The gradient ascent method used for training the loss functions. iv) The activation functions used. For DCAN, we use $tanh$ for the hidden layers and $sigmoid$ for the output. We set the clustering iteration to one in Table 2 and a sufficient statistics of at least 600 points in the raw latent space to estimate the cluster parameters. The number of cluster is the same as the number of classes. We use a minibatch size of $N=16$ samples, 500 iterations and we use gradient ascent with momentum with a learning rate between $10^{-3}$ and $10^{-5}$. We use accuracy (ACC) [15] to evaluate the performance of our clustering. MNIST: A well cited 10 digit classes dataset with no background but it has more samples than USPS. The train and test sample size is at 50K and 10K. Our settings are $196-384-256$ for the encoder and $256-16-1$ for the discriminator. The input dimension is originally at $28\times 28$ but we have downsampled it to $14\times 14$ and flattened to 196. In Table 2, DCAN using raw pixel information obtained $0.8565$ for ACC which is on par or better than deep clustering methods including DEC, ABC and DC-GMM. However on MNIST, most deep clustering methods could not obtain the performance of recent deep learning methods such as KINGDRA-LADDER-1M (a.k.a. KINGDRA) and IMSAT as their goals are much more ACC result oriented. For KINGDRA, it combines several ideas together including ladder network [16] and psuedo-label learning [17] for the neural network and using ensemble clustering. In DAC, their approach is based on pairwise classification. In IMSAT-RPT, the neural network was trained by augmenting the training dataset. In our case, we simplified our focused on deep clustering using adversarial net. DCAN penalizes samples for being an outlier from their cluster centers and does so uniquely by using the discriminator to penalize the encoder. The original objective of deep clustering is specifically aimed at clustering. In some datasets, the latent space of this objective do not necessarily co-serve as the most important factor for raw pixel feature extraction. Thus, we also used ResNet18 pretrained model as a feature extractor for MNIST, prior to deep clustering. This simple step allows the ACC for DCAN to improved to 0.995 (not shown in Table 2) and in fact outperforms KINGDRA. Reuters-10k: The Reuters-10k dataset according to [18, 19] contains 10K samples with 4 classes and the feature dimension is 2000. End-to-end learning is performed on this dataset. We compared DCAN to other unsupervised deep learning approaches. DCAN uses $2000-100-500$ and $500-100-1$ respectively for the encoder and discriminator. In Table 2, we observed that Kmeans alone can achieve an ACC of 0.6018 on the raw feature dimension. Most deep learning methods can achieve between 0.69 to 0.73 on this dataset. DCAN was able to obtain the best ACC at 0.7867. CIFAR10: A 10 classes object categorization dataset with 50K training and 10K testing of real images. On CIFAR10, most methods have difficulty performing end-to-end learning. As compared to deep ConvNet which utilizes convolutional neural network and supervised learning for a specific task (i.e. feature extraction), deep clustering typically rely on fewer hidden layer in the encoder for end-to-end learning (i.e. both feature extraction and clustering). Thus, we followed the experiment setup in KINGDRA [20] by using the “avg pool” layer of ResNet50 [21] (ImageNet pretrained) as a feature extractor with a dimension of 2048. In Table 2, DCAN uses $2048-1024-512-128$ and $128-16-1$ respectively for the encoder and discriminator. Despite the stronger performance of DAC, IMSAT and KINGDRA on MNIST, we outperformed these comparisons on CIFAR10 at $0.5844$. This is despite the fact that both KINGDRA and IMSAT also use ResNet50 (ImageNet pretrained) for CIFAR10.
Approach MNIST Reuters10k CIFAR10
ABC [7] 0.760 0.7019 0.435
DEC [19] 0.843 0.7217 0.469
DC-GMM [18] 0.8555 0.6906 -
AAE [13] 0.8348 0.6982 -
IMSAT-RPT [22] 0.896 0.719 0.455
KINGDRA [20] 0.985 0.705 0.546
DCAN (proposed) 0.8565 0.7867 0.5844
Table 2: ACC Benchmark
4 Conclusion
The objective of deep clustering is to optimize the autoencoder latent space with clustering information. The key to this technique lies in a loss function that minimizes both clustering and encoder in the latent space. Despite the recent success of deep clustering, there are some concerns: i) How to use probabilistic approach for the loss function? ii) Can we backpropagate the neural network when there is no closed form solution for the probabilistic function? A JS divergence approach does not suffer from asymmetry like KL divergence loss. Unlike the latter, there is no closed form solution to the JS divergence when backpropagating the network. This paper addresses the above two concerns by proposing a novel deep clustering approach based on adversarial net to estimate JS divergence for deep clustering.